How many electrons are present in an atom with its 1s, 2s, 2p, 3s, 3p, 3d, 4s, 4p subshells filled?
Answer #1

Given that ,

1s, 2s, 2p, 3s, 3p , 3d, 4s, 4p

Subshell: it is the subdivision of shell which is denoted by s: sharp, p: principle,D : diffusion and f is fundamental etc.

Shell Subshell Total Number of Electrons in Shell
1st Shell 1s 2
2nd Shell 2s, 2p 2 + 6 = 8
3rd Shell 3s, 3p, 3d 2 + 6 + 10 = 18
4th Shell 4s, 4p, 4d, 4f 2 + 6 + 10 + 14 = 32

Number of electron s subshell is 2e

P subshell is 6e

d subshell is 10e

Total number of electron present in electronic configuration

2+2+6+2+6+10+2+6=36